tickets cost $6.00 for adults and $3.00 for students.total of 1050 tickets worth $4200 sold.How many adult tickets sold.
Answer by mananth(16946)

x Adult tickets
y Student tickets
Total tickets sold
1.00 x + 1.00 y = 1,050.00 .............1
Value of tickets
6.00 x + 3.00 y = 4,200.00 .............2
Eliminate y
multiply (1)by -3.00
Multiply (2) by 1.00
-3.00 x \ -3.00 y = -3,150.00
6.00 x + 3.00 y = 4,200.00
Add the two equations
3.00 x = 1,050.00
/ 3.00
x = 350.00
plug value of x in (1)
1.00 x + 1.00 y = 1,050.00
350.00 + y = 1,050.00
y = 1,050.00-350
y = 700.00
y = 700.00
350.00 Adult tickets
700.00 Student tickets
m.ananth@hotmail.ca